# Knowledge Augmentation in Federation: Rethinking What Collaborative Learning Can Bring Back to Decentralized Data

### v2

@@ VI-C Summary
In this section we highlight emerging studies that work on the dimensions in which data are represented through collaborative effort from multiple parties. We argue that they symbolize an important approach to knowledge augmentation in federation. On the one hand, federated data correction attempts to amend the misinformation carried by the data to avoid erroneous knowledge, with particular focus on the labels. Reference knowledge, in most of the label-correction methods, is encapsulated in the global model. On the other hand, federated feature engineering focuses on identifying the most important attributes or the most expressive combination of them. In this context, the statistical properties of features are often utilized as the measure of knowledge they can provide.
Although different in methodologies (summarized in Table IV, the two lines of research have some common points in terms of implementation. First, both data correction and feature engineering are typically implemented as part of a multi-stage procedure in the federation. For example, data correction is usually cooperated by the clients in the pre-processing stage before the actual training process such as FL. Similarly, federated feature engineering usually involves repetitive local feature assessment and client-server information exchange. In addition, we discover that both approaches can be sensitive to the discrepancy of data distribution across different domains. A unified posterior (that describes the conditional probability of labels) or an optimal feature subspace may not work well for every individual client in the federated system.
@@ VII Open Challenges and Questions
Through revisiting diverse approaches towards more effective, efficient learning from various perspectives of knowledge, we argue that there is a large space ahead to explore in the domain. The advance in knowledge discovery combined with the promising potential of federated learning is expected to open up new opportunities for knowledge-centric, learning-driven collaboration in the future. In the paper we highlight urgent challenges and open questions that involve but are not limited to the ways knowledge is learned, represented and exchanged as well as how to enhance Large Language Models (LLMs) with KAF. We expect to offer unique insights for research beyond knowledge for learning.
VII-A Decentralized Peer-to-Peer Knowledge Augmentation
Extensive effort has been made in decentralized machine learning since it has natural advantages in scalability, flexibility and byzantine-robustness [111, 112]. However, existing approaches are mostly learning-centric, which means that their objectives are shaped towards the optimization of a specific group of models or a family of relevant tasks. This makes the collaboration costly and hard to reproduce, because the model as the knowledge carrier is very sensitive to the uncertainty in federated training. This uncertainty boils down to the population involved as well as the sequence of knowledge propagation in the network, which also raises the cost of reproducing the model.
Recent studies on decentralized matrix factorization (DMF) present a domain-specific path to expand local knowledge domain via peer-to-peer (P2P) communications [113, 114, 115]. DMF can be regarded as a special case of decentralized knowledge augmentation. With a novel vision, Abdelmoniem et al. [116] described a decentralized knowledge routing framework where a client can describe the knowledge requested and find another well-matched client for collaborative learning. This is done by sharing artificial samples with the proxy devices (called knowledge discoverers) responsible for request routing and handling. Even though the objective employed is still task-specific and learning-oriented (deficiencies discussed in Sec. I still apply), it inspires us to explore new directions for decentralized knowledge exchange and knowledge matching between domains.
Clues for further study: knowledge abstraction and matching, blockchain, zero-knowledge proof, and knowledge routing.
@@ VII-B Light-weight Knowledge Courier and Efficient Exchange
A key factor in knowledge augmentation in federation is the exchange of knowledge between distributed local domains. For example, the item factor matrix is shared in the federation as the carrier of domain-specific knowledge in federated collaborative filtering, whilst the global model takes this role in federated label denoising. Whatever form it takes for knowledge exchange, the reduction of communication cost persists as a top priority in practice.
@@
On the application layer, the roadmap for efficient knowledge exchange mainly encompasses two approaches: i) decreasing the communication payload incurred by the courier of knowledge, and ii) reducing the number of communication rounds if iterative interactions are needed. Although FL is more communication-efficient than traditional distributed learning, it is worth noting that models, model updates and gradients are “heavy”. Provided that it is not necessary to exchange the “full picture” of knowledge in many situations such as federated knowledge filtering and label space correction, how to craft a flexible protocol to avoid information redundancy emerges as new problem to be resolved.
@@
Clues for further study: model-free knowledge representation, information bottleneck, and Bayesian compression.
VII-C Federated LLM Augmentation with Exogenous Knowledge
It is impossible for any model, even as complex as the latest LLMs [5, 6, 117], to learn and store comprehensive knowledge that covers all domains of training data. Moreover, language models usually suffer in reproducing accurate knowledge, a.k.a. the problem of hallucinations. Recent studies have paid attention to the federated Retrieval Augmentation Generation (RAG) [118], which is a promising technique for introducing exogenous knowledge to assist in model reasoning. These attempts target at addressing the “information scattering” effect [119] by searching multiple domains, yet the trade-off between performance gain and retrieval cost is mostly overlooked. The effectiveness and efficiency of federated RAG hinge on two factors, that is, whether a selected exogenous knowledge base well matches the given query, and whether the content retrieved can provide qualified knowledge. On the one hand, it is very expensive in time to reliably determine whether a huge database contains the piece of knowledge desired. On the other hand, the scarcity of user feedback brings difficulties for assessing the usefulness of contents retrieved. To harness the potential of federated retrieval for enhanced LLM services, deeper investigation into the evaluation of exogenous knowledge is needed.
@@ VII-C Federated LLM Augmentation with Exogenous Knowledge
Clues for further study: federated RAG, federated search systems, query-content evaluation, and knowledge base summary.
VII-D KAF for Unlabeled data
We observed that the majority of researches in relevance to knowledge augmentation and federated learning were designed for or evaluated on supervised learning tasks. This is mainly attributed to the prior knowledge implied by the labels that usually serve as the most important criterion for many operations. For example, feature engineering typically relies on feature-label mutual information, and data re-balancing strategies re-weight data points by class label. It is also worth mentioning that most existing empirical studies on federated learning associate non-IIDness to the marginal distribution of labels. However, this largely limits their application to a broader categories of unlabeled data such as text corpora, surveillance footage and browsing records. We argue that the main challenge for KAF over unlabeled data lies in the perplexity of knowledge within each domain. This means that, without the a priori from annotations, the increased uncertainty of data (e.g., measured in entropy) may result in lower confidence in the knowledge learned. In this case, more attention should be paid to the intrinsic structures of data and the impact of non-IID data beyond classification.
Clues for further study: active learning, contrastive learning, soft labels, and knowledge distillation.
VII-E Can KAF Make Independent Training Superior over FL?
@@
We argue that refined representation of knowledge is what collaborative learning can bring back to data. This is the key point that distinguishes KAF from what have been explored in the relevant field such as Federated Learning. As discussed, knowledge augmentation in federation can enhance the local knowledge in each domain by means of expansion, filtering and correction. This creates possibility for each participant to learn independently after the augmentation. But how competitive will it be as compared to Federated Learning? The answer is uncertain and hinges on multiple factors including local data size, the test data distribution, and the utility of knowledge gained via KAF. In terms of performance, KAF can increase the size and diversity of samples necessary for stochastic gradient-based local training, with augmented knowledge from foreign domains for generalizable model. In terms of cost, FL struggles in communication efficiency with large models and huge client population, whilst KAF plus independent training has tempting potential when equipped with efficient knowledge exchange protocols. From both perspectives, the superiority of KAF is still an open-ended question.
Clues for further study: benchmarking, sample-efficient learning, and marginal gains of collaboration.
VII-F Can Knowledge be Borrowed and Unlearned in a Federation?
“The right to be forgotten”, as stated in privacy-protecting regulations, has given rise to the interest in the research of Machine Unlearning [120, 121]. With potential value in achieving security, privacy, usability, and fidelity in various application scenarios [121], the concept of machine unlearning has also been adopted in FL systems to enable a specific client’s knowledge to be removed from what has been learned by the model [122]. However, existing approaches are basically limited to the unlearning of class-oriented knowledge [123]. More importantly, the process of unlearning can cost as much as re-training in terms of time and resources. This calls for further study on approximate federated unlearning to enable efficient and generalizable knowledge erasure. In the context of KAF, collaboratively augmented data can be reckoned as “borrowed” knowledge. New mechanisms are required in order to accurately identify the synthetic data points that characterize the designated part of knowledge to be forgotten.
Clues for further study: federated machine unlearning, model inversion, gated knowledge transfer, noise learning, and influence-aware filtration.
@@ VIII Conclusion
As the value of data assets and privacy issues arise, it is of utmost importance for organizations to make full use of proprietary data whilst seeking cooperation with their counterparts. Although the collaborative paradigm of federated learning has paved the way for the sharing of knowledge across domains with privacy preservation, research in the field hardly explored the possibilities beyond model-centric learning. We argue that knowledge sharing is not just limited to the purposes of model optimization, but can also serve as the key to local domain augmentation, which substantially diversifies the way how participants can benefit from the federation. With this vision, we in this paper present knowledge augmentation in federation, a conceptual paradigm for the expansion, filtration and correction of knowledge across decentralized domains. We revisit emerging research in the field via the lens of knowledge and provide a category of studies according to their methodologies. Through comparison and discussion we point out that the carrier of knowledge and the protocol for knowledge exchange are the most influential factors for KAF to be effective and efficient. Based on our investigation, we further highlight a number of open challenges that are not adequately addressed by existing effort, and also pose two open-ended questions to inspire more theoretical and empirical exploration in the relevant threads of research.
Acknowledgment
This work is supported by the National Natural Science Foundation of China, National Social Science Foundation of China Post-funded project, Basic and Applied Basic Research Project of Guangzhou, and the Fundamental Research Funds for the Central Universities.

### v3

@@ VI-C Summary
In this section we highlight emerging studies that tailor the dimensions in which data are represented. Federated data correction attempts to amend the misinformation carried by the data to avoid erroneous knowledge, with particular focus on the labels. Reference knowledge, in most of the label-correction methods, is encapsulated in a global model. This limits their application to task-agnostic scenarios. Federated feature engineering focuses on identifying the most important attributes or the most expressive combination of them. In this context, the statistical properties of features are often utilized as the measure of knowledge they can provide. Their pros and cons are summarized in Table IV.
Although being different in methodologies, the two lines of research have much in common. First, both data correction and feature engineering are typically implemented as part of a multi-stage procedure in the federation. For example, data correction is usually cooperated by the participants in the pre-processing stage before the downstream tasks. Similarly, federated feature engineering usually involves repetitive local feature assessment and client-server information exchange. In addition, we discover that both approaches can be sensitive to the discrepancy of data distribution across different domains. A unified posterior (that describes the conditional probability of labels) or an optimal feature subspace may not work well for every individual participant in the federated system.
@@ VII Open Challenges and Questions
Through revisiting diverse approaches which bring generic knowledge into local data domains in a federated system, we argue that there is a large space ahead to explore especially in the aspects of efficiency and efficacy. In this section we highlight opportunities for future research in relevance to TAF.
VII-A Decentralized Peer-to-Peer Knowledge Seeking
Decentralization has natural advantages in scalability, flexibility and byzantine-robustness [vincent2023systematic]. Recent studies on decentralized matrix factorization (DMF) present a domain-specific path to expand local data via peer-to-peer (P2P) communications [hou2023bfrecsys, huang2024towards]. DMF can be regarded as a special case of decentralized data augmentation, yet it is hard to deal with general forms of data other than matrices. With a novel vision, Abdelmoniem et al. [abdelmoniem2025] described a decentralized knowledge routing framework where a participant can describe the knowledge requested and find another well-matched peer for collaborative learning. This is done by sharing artificial samples with the proxy devices (called knowledge discoverers) responsible for request routing and handling. Though the objective employed is still task-specific and learning-oriented, it inspires us to explore new directions for decentralized knowledge exchange and knowledge matching between domains.
Clues for further study: knowledge abstraction and matching [wu2023fedprof, ahmed2024fi], blockchain, zero-knowledge proof, and knowledge routing [abdelmoniem2025].
@@ VII-B Light-weight Knowledge Courier and Efficient Exchange
A key factor in task-agnostic federation is the exchange of knowledge between distributed local domains. For example, the item factor matrix is shared as the carrier of domain-specific knowledge in federated collaborative filtering, whilst the global model takes this role in federated label denoising. Whatever form it takes for knowledge exchange, the reduction of communication cost persists as a top priority in practice.
@@
On the application layer, the roadmap for efficient knowledge exchange mainly encompasses two approaches: i) decreasing the communication payload incurred by the courier of knowledge, and ii) reducing the number of communication rounds if iterative interactions are needed. Existing studies often exchange models, but it is worth noting that models, model updates and gradients are “heavy”. Provided that it is not necessary to exchange the “full picture” of knowledge in many situations such as federated data filtering and harmonization, how to craft a flexible protocol to avoid information redundancy emerges as new problem to be resolved.
@@
Clues for further study: model-free knowledge representation, information bottleneck [yang2024ib], and Bayesian compression [yang2025continual].
VII-C Federated Retrieval of Exogenous Knowledge with Confidentiality
Latest LLM-empowered dialogue systems [guo2025deepseek, openaio1systemcard] commonly adopt Retrieval-Augmented Generation (RAG) to bring in up-to-date, verified information so as to produce accurate knowledge, a.k.a. reducing hallucinations. With rising awareness of intellectual properties, recent studies have paid attention to the federated Retrieval-Augmented Generation (FedRAG) [shojaee2025federated] with requirements of context confidentiality [addison2024cfedrag]. This makes a special case of TAF where knowledge exchange (including queries and context) happens between multiple knowledge bases (KBs) and a reasoning agent. In this scenario, few studies stand on the KB side and they hardly account for the cost of communication and the price of retrieval service. On the one hand, it is very expensive in time to reliably determine whether a huge database contains the piece of knowledge desired. On the other hand, the scarcity of user feedback brings difficulties for assessing the usefulness of contents retrieved. In addition, it is worthwhile to explore how multiple knowledge bases can cooperate with each other, e.g., in a TAF setting, to provide refined and logically consolidated retrieval capability.
@@ VII-C Federated Retrieval of Exogenous Knowledge with Confidentiality
Clues for further study: federated RAG with confidentiality [addison2024cfedrag], federated search systems [guerraoui2025efficient], query-content evaluation [yu2024knowledge], and scattered knowledge bases [wu2025multirag].
VII-D Task-agnostic Collaboration over Incomplete Data
We observed that the majority of researches in the federated setting assume the completeness of data, overlooking the fact that real-world data have missing attributes and labels. Especially, many existing studies rely on data labels for knowledge extraction and empirically associate non-IIDness to the marginal distribution of labels. For example, feature engineering typically relies on feature-label mutual information, and data re-balancing strategies re-weight data points by class label. This means that they cannot apply to collaboration over incomplete data. This largely limits their application to a broader categories of unlabeled data such as text corpora, surveillance footage and browsing records. We argue that it is important while challenging for the data owners in TAF to represent incomplete data in a traditional form. For example, it is impossible to swap class-wise information over unlabeled datasets. In this case, we suggest further exploration into flexible representations of data based on their intrinsic structures and robust protocols for knowledge exchange between parties of varied data conditions.
Clues for further study: fuzzy representations, auxiliary supervision [yang2024depth], sample fusion, and knowledge-data alignment [zhao2024weakly] .
VII-E Can TAF Make Autonomy Superior over FL?
@@
We argue that the gain of collaborative learning becomes marginal when the participants’ data get larger and diversified. This is the key factor that limits the sustainability of learning-centric collaboration such as FL. On the contrary, task-agnostic federation liberates the participants from having to work on a specific model. This creates possibility for each participant to learn independently after the collaboration. But how competitive will it be regarding the downstream task when compared to federated learning? The answer is uncertain and hinges on multiple factors including local data size, the test data distribution, and the utility of data gained via TAF. In terms of performance, TAF can increase the size and diversity of samples necessary for independent training, with augmented knowledge from foreign domains for generalizable model. In terms of cost, FL struggles in communication efficiency with large models and huge participant population, whilst TAF plus independent training has tempting potential when equipped with efficient knowledge exchange protocols. From both perspectives, it is still an open-ended question.
Clues for further study: benchmarking, query-based knowledge transfer [alballa2025qkt], and autonomous organization [zhang2025srfl].
VII-F Can TAF Support Knowledge Erasure?
“The right to be forgotten”, as stated in privacy-protecting regulations, has motivated recent studies in unlearning knowledge gained in a group [chen2024fmu]. However, many existing unlearning approaches sit on unrealistic assumptions such as storing the entire trace of gradients or forcing all participants to engage in the unlearning stage [tarun2024fast]. More importantly, they are yet to be evaluated on realistic unlearning scenarios where the shared knowledge needs to be “withdrawn” any time on demand. In contrast to these traditional settings, TAF opens up new opportunities for knowledge erasure. On the one hand, collaboratively augmented data can be reckoned as “borrowed” knowledge which is easy to remove. On the other hand, we believe it is also possible to accelerate any downstream unlearning task by exchanging artificial samples dedicated for “overwriting” specific knowledge. In both ways TAF can support efficient knowledge erasure. It also calls for innovative data engineering and knowledge exchange protocols in response to this special demand.
Clues for further study: Augmentation-based unlearning [falcao2025data, Choi2024CVPR], Adversarial Mixup [peng2025adversarialmixup], inhibited synthetic data [zhang2025toward], influence-aware filtration [dang2025efficient], and noise learning [chien2024langevin].
@@ VIII Conclusion
As the value of data assets and privacy issues arise, it is of utmost importance for data owners to make full use of proprietary information whilst seeking cooperation with their counterparts. Although federated learning-based paradigms have paved the way for privacy-preserving collaboration, its learning-centric nature greatly limits its feasibility over cost-sensitive, self-interested data owners. We argue that it is worth thinking of a new scenario by shifting the focus to data-centric and task-irrelevant cooperation, which substantially diversifies the way how participants can benefit from the federation. With this vision, we first conceptualize a new scenario of collaboration termed Task-Agnostic Federation, and then investigate latest research that enables exchange of generic knowledge to achieve collaborative data expansion, data refinement and data harmonization over decentralized data. Through a comprehensive literature review, we point out that the carrier of knowledge and the protocol for knowledge exchange are the most influential factors for TAF to be effective, efficient and extendable. Based on our investigation, we further highlight a number of open challenges that are not yet adequately addressed by existing effort, and also pose two open-ended questions to inspire more theoretical and empirical exploration in the domain.
Acknowledgement
This work is supported by the National Natural Science Foundation of China (62402198, U23B2027), National Social Science Foundation of China Post-funded Project (20FGLB034), Basic and Applied Basic Research Project of Guangzhou (2025A04J2212), the Fundamental Research Funds for the Central Universities (21624348), and New Generation Artificial Intelligence-National Science and Technology Major Project (2025ZD0123605).
An earlier preprint version of this article entitled “Knowledge Augmentation in Federation: Rethinking What Collaborative Learning Can Bring Back to Decentralized Data” was posted on arXiv (DOI arXiv:2503.03140).